OpenStudy (anonymous):

Which table best represents the steps to solve the equation 8 + x over 3 = 4?

OpenStudy (anonymous):

\[\frac{8+x}{3}=4\\ 8+x=4\times 3\\ 8+x=12\\ x=4\] is that choice?

OpenStudy (anonymous):

no